OpenStudy (anonymous):

What is the solution of the equationlog 2 (5x−2)=3

OpenStudy (unklerhaukus):

is that \[\log_2(5x-2)=3\]?

OpenStudy (unklerhaukus):

the first step is to raise both sides of the equation to the power of two \[\large2^{\log_2(5x-2)}=2^3\] the left hand side will simplify \[\color{red}{\cancel{\large2}^{\log_2\color{black}{(5x-2)}}}=2^3\]\[5x-2=2^3\]